lisa leaves philadelphia at 9:00am for Boston at an average speed of 45 mph. Two hours later, Mike leaves Philadelphia for Washington, DC traveling at an average speed of 60 mph. At what time will Lisa and Mike be 300 miles apart?
Answer by lwsshak3(11628)

lisa leaves philadelphia at 9:00am for Boston at an average speed of 45 mph. Two hours later, Mike leaves Philadelphia for Washington, DC traveling at an average speed of 60 mph. At what time will Lisa and Mike be 300 miles apart?
***
let x=mike's travel time
x+2=lisa's travel time
distance=speed*travel time
..
45(x+2)+60x=300
45x+90+60x=300
105x=210
x≈2
x+2=4
At what time will Lisa and Mike be 300 miles apart? 9+4hrs=1:00 pm
***